An object that weighs 1000 N floats with 80% of its volume submerged below water. What is the boyance force? What is the volume of the object?
Answer #1

boyance force= weight of the object
boyance force= 1000 N  

density of object = 1000x80/100 = 8000 kg/m3

mass of object = 1000/9.81 = 102 kg  

volume = mass/density  
= 0.127 m3